Hello, This is Janice again.
To: venugopalramana(480)
You recently helped me with a math problem. But I was confused about something where did you get the 111mg from?

Answer by venugopalramana(3286)

SEE MY ANSWER
QUOTE:
FIRST FIND MILI GRAM EQUIVALENT MOLECULAR WEIGHT OF CALCIUM CHLORIDE FROM ITS FORMULA ..THERE IS 1 ATOM OF CALCIUM AND 2 ATOMS OF CHLORINE IN IT.HENCE M.WT. =1*CA+2*CL=1*40+2*35.5=111...
UNQUOTE:
MOLECULAR WEIGHT IS THE SUM OF ALL COMPONENT'S WEIGHT IN 1 MOLECULE OF THE COMPOUND.IN CALCIUM CHLORIDE COMPOUND ,AS PER THE FORMULA CaCl2,THERE IS 1 ATOM OF CALCIUM(AS SHOWN BY Ca OR Ca1 IN EFFECT) AND 2 ATOMS OF CHLORINE(AS SHOWN BY Cl2).ATOMIC WEIGHT OF CALCIUM IS GIVEN AS 40 AND THAT OF CHLORINE IS GIVEN AS 35.5.SO WE MULTIPLIED 40 WITH 1 FOR 1 ATOM OF CALCIUM AND 35.5 WITH 2 FOR 2 ATOMS OF CHLORINE IN THE COMPOUND AND ADDED THEM UP TO GET THE TOTAL MOLECULAR WEIGHT AS 111 AS SHOWN ABOVE.
NOW A WORD ABOUT CALLING THESE AS WEGHTS IN GRAMS OR MILLIGRAMS ETC.STRICTLY SPEAKING THEY ARE GRAM OR MILLIGRAM EQUIVALENTS.THAT IS IF WE CONSIDER 111 AS MILLIGRAMS OF CALCIUM CHLORIDE,THEN IT IS CALLED MILLIGRAM MOLECULAR WEIGHT AND ALL COMPONENTS OF IT LIKE CALCIUM AND CHLORINE ARE ALSO TAKEN AS MILLIGRAMS.IF TH MOLECULR WEIGHT IS TKEN AS GRAMS OR KILOGRAMS ,THEN SAME MEASURE IS TAKEN FOR ITS COMPONENTS.HOWEVER ,THE ACTUAL WEIGHT OF 1 MOLECULE IS FAR FAR FAR FAR LESS THAN A MILLI...MILLI..GRAM,WHICH IS DIFFICULT AND TOO SMALL FOR USE IN OUR APPLICTIONS.HENCE WE USE THIS TERM OF EQUIVALENT MILLIGRAMS OR GRAMS ETC..AND USE IT. AS LONG AS WE ARE TALKING OF RATIOS IT DOES NOT MATTER.
VENUGOPAL
SO IN 111 MG OF CACL2 WE HAVE 40 MG.EQV. OF CALCIUM
IN 250 MG OF CACL2 WE SHALL HAVE 250*40/111=90 MG.APPROXIMATELY
- Show quoted text -